### v1

# Utilizing the Janus MoSSe surface polarization property in complementary metal–oxide–semiconductor field effect transistor design

Abstract
Janus transition metal dichalcogenides (JTMDs) have attracted much attention because of their outstanding electronic and optical properties. The additional out-of-plane dipole in JTMDs can form n- and p-like Ohmic contacts, and this may be used in device applications such as pin diodes and photovoltaic cells. In this study, we exploit this property to design n- and p-type metal–oxide–semiconductor field effect transistors (MOSFETs). First, we use density-functional theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe. The intrinsic dipole of MoSSe causes band bending at both the metal/MoSSe and MoSSe/metal interfaces, resulting in electron and hole accumulation to form n- and p-type Ohmic contact regions. We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, on the basis of which we design complementary MOSFETs. Our results demonstrate that JTMDs can be used to make n- and p-MOSFETs in the same layer without the need for any extra doping.
I Introduction
In the field of two-dimensional (2D) semiconducting materials, research on transition metal dichalcogenides (TMDs) has been flourishing in recent years owing to their remarkable properties, such as relatively high carrier mobilities,Chhowalla, Jena, and Zhang (2016a); Chen et al. (2022a); Radisavljevic et al. (2011) suitable bandgap energies,Ye, Zhang, and Yap (2017); Chaves et al. (2020) and an atomic thickness that is modifiable.Serna et al. (2016); Zhang et al. (2013); Orofeo et al. (2014) These properties make 2D TMDs suitable for various device applicationsLiu et al. (2021); Cho, Park, and Kwon (2015); Ying-Chuan_KMC; Xiong et al. (2021); 2D_Photo, especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs). Liu et al. (2021); Chhowalla, Jena, and Zhang (2016b); 2D_FET_Pin-Fang These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials. For example, MoS and WS are two commonly selected materials for n-type channels, and -MoTe and WSe are both good candidates for p-type channels.Radisavljevic et al. (2011); Pezeshki et al. (2016); Ovchinnikov et al. (2014); Cho, Park, and Kwon (2015); Zhang et al. (2021)
It is well known that the contact characteristics of a FET device will strongly affect its – characteristics and performance. For low contact resistance, Ohmic contacts are usually required at the semiconductor–metal interface to enhance the electrical performance of FETs while maintaining a linear – curve.Sze, Li, and Ng (2021); Allain et al. (2015); Zheng et al. (2021) Unfortunately, Schottky barriers (SBs) are usually formed for most TMD/metal contacts.Allain et al. (2015); Kang et al. (2014); Ouyang, Xiong, and Jing (2018); Tang et al. (2019); Zheng et al. (2021) For an FET with SB contacts, carriers must overcome the energy barrier to transport from metal into the TMD channel region, impairing the device function by lowering the injection efficiency.Tung (2014) A typical technique used to reduce contact resistance in modern semiconductor technology is substitutional doping.Younas, Zhou, and Hinkle (2023) However, there are better ways to apply such a doping method to 2D materials, since their characteristics will be heavily altered after doping. Additionally, the absence of dangling bonds on the surface of a 2D material makes solid interface connections with metal challenging, resulting in larger contact resistance.Allain et al. (2015) Hence, most of the effort required when engineering TMD/metal contacts is associated with finding a suitable metal to adjust the work function. For n-FETs, a material with a smaller work function is needed, whereas a metal with a high work function is often selected for p-FETs. Several studies have shown that TMD/metal heterostructures such as MoS/Cu, WS/Cu, and WSe/Ti could provide n-type SB contacts,Kang et al. (2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe/Pd and WSe/Au.Kang et al. (2014); Kong et al. (2020) However, only a specific TMD/metal combination is acceptable for Ohmic contact formation, making TMD-based FET contact design more challenging.Zhao and Schwingenschlögl (2022)
To combine n- and p-type FETs into a complementary metal–oxide–semiconductor FET (CMOSFET) device, horizontally combing two different TMD channels to form a CMOSFET is a common approach.Liu et al. (2014); Das, Dubey, and Roelofs (2014); Cho, Park, and Kwon (2015); Xiong et al. (2021) With regard to fabrication, however, this CMOSFET design with two different materials has the disadvantage that the two materials need to be transferred onto a single substrate, with the risk of introducing defects or contamination.Ullah et al. (2021) Furthermore, two contact metals need to be considered separately, resulting in higher processing complexity and a more time-consuming fabrication procedure. Recently, ambipolar TMDs such as MoTeLin et al. (2014) and ReSeLee et al. (2019) have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
The use of a novel type of materials, namely, Janus transition metal dichalcogenides (JTMDs), MXY, where M transition metal and X, Y S, Se, or Te, with X Y, is another possible way to fabricate a CMOSFET using a single material. In recent years, several JTMDs have been successfully synthesized via different techniques. Monolayer MoSSe, for instance, was synthesized utilizing chemical vapor deposition combined with plasma strippingLu et al. (2017) or a selective epitaxy atomic replacement process,Trivedi et al. (2020) while Lin et al. Lin et al. (2020) prepared Janus WSSe by pulsed laser deposition. Previous theoretical studies of these JTMDs have provided unmistakable proof that their phonon spectra are stable and devoid of imaginary frequencies, verifying their fabrication stability.Petrić et al. (2021); Chaurasiya, Gupta, and Dixit (2019) Compared with TMDs, JTMDs have two different chalcogen-atom layers, which create an asymmetric sandwich configuration that provides an out-of-plane electric field. This out-of-plane field has the characteristic of a dipole, which enables potential applications in piezoelectric, spintronics, catalysts, and gas sensing devices.Zhang et al. (2022, 2020); Kong et al. (2023) Also, simulation studies have shown that this intrinsic dipole feature of JTMDs can enable the formation of both n- and p-type contacts simultaneously without the need for additional doping when different sides are attached to graphene,Chen et al. (2022b); Palsgaard et al. (2018); Liu et al. (2020) and this can be incorporated into the design of pin-diodePalsgaard et al. (2018) or photovoltaicLiu et al. (2020) devices. Moreover, theoretical studies have shown that the SB at contact could be substantially reduced when JTMD MoSSe are attached to metals,Zhao and Schwingenschlögl (2022) owing to the intrinsic dipole. The above-mentioned JTMD/electrode contact systems may even attain Ohmic contact under certain conditions.Chen et al. (2022b); Zhao and Schwingenschlögl (2022) However, unlike optoelectronic devices, the exploitation of the advantages of n- and p-type Ohmic contacts in FET applications has not yet been explored. Furthermore, there has not been much work on utilizing this intrinsic dipole property to form FETs or CMOSFETs without doping.
In this study, we propose a CMOSFET design concept utilizing the intrinsic dipole of a JTMD. The material MoSSe is selected for investigation, since it was the first JTMD and is the most commonly synthesized.Lu et al. (2017); Trivedi et al. (2020) The remainder of this paper is organized as follows. First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, bilayer MoSSe (2L-MoSSe)/graphene and MoSSe/Al (111) face and MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT). Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson–drift–diffusion solver (2D-DDCC) developed in our laboratory.Wu
II Computational details
II.1 Heterostructure and layered MoSSe DFT simulations
The heterostructure calculations are performed with Quantum ESPRESSO,Giannozzi et al. (2009) an open-source DFT calculation software package employing the generalized gradient approximation (GGA) of the exchange-correlation functional and norm-conserving VanderbiltHamann, Schlüter, and Chiang (1979) pseudopotentials combined with DFT-D3 van der Waals correction.Grimme et al. (2010) An Monkhorst–Pack sampling is used. The geometry is refined until the maximal residual forces are less than 0.015 eV/Å, at which point MoSSe heterostructures are subjected to the dipole correction to negate the artificial electric field caused by their asymmetric structures. The vacuum region is set to 20 Å to prevent interactions between nearby supercells. For MoSSe contact simulation, we attach a 1 supercell of graphene to a two-layer AA stacking supercell of MoSSe, A supercell of six layers of (111) Al to a 1 supercell of MoSSe, and a 1 supercell of six layers of (111) Cu to a supercell of MoSSe. A simulation model is created for each metal’s energetically advantageous surface,Skriver and Rosengaard (1992) and four outward metal layers are fixed in the relaxation process. For layered MoSSe, we construct an AAA stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here. The only difference is that the mesh is set to in the MoSSe band structure calculation.
II.2 Device simulation
The 2D-DDCC solver is used to examine the device performance. This solver was developed in our laboratory and has been used on many different kinds of devices. More detailed information can be found on our website.Wu Both the Poisson and drift–diffusion equations are solved self-consistently by the solver. The equations are as follows:
where is the dielectric constant, is the electric potential, is the acceptor’s doping concentration, is the donor’s doping concentration, is the electron concentration, is the hole concentration, and are the current densities of electrons and holes, respectively, and and are the electron and hole diffusion coefficients, respectively. This solver serves as the basis for all of the device simulations discussed in this article, with the exception of the MoSSe channel region calculations.
III Results and discussion
To study the interactions between MoSSe and contacts, we construct the S and Se sides of the MoSSe heterostructure with graphene and two common electrode metals (namely, Al and Cu). Figure 1(a) shows the calculated heterostructure systems that can form low SB or Ohmic contacts. The top views are shown in Figs. 1(b) and 1(c), where the lattice mismatches are 1.11%, 0.44%, and 2.97% for the 2L-MoSSe/graphene, MoSSe/Al, and MoSSe/Cu systems, respectively. Figures 1(e)–1(j) show the projected band structures, and Figs. 1(e)–1(g) representing the S-face contact of MoSSe with the electrode material, and Figs. 1(h)–1(j) representing the Se-face attachment. As can be observed, if graphene contacts a two-layer MoSSe material on the Se side [Fig. 1(e)], then the Fermi level of the entire system is 0.1 eV lower than the observable conduction band of the MoSSe material, suggesting an n-type Schottky barrier. On the other hand, if graphene contacts a two-layer MoSSe material on the S side [Fig. 1(h)], then we see a p-type Ohmic contact. These results are consistent with the results of previous JTMD contact studies.Palsgaard et al. (2018); Chen et al. (2022b) According to our DFT calculation, monolayer MoSSe has a strong V/cm intrinsic dipole field in the out-of-plane direction, which may be the reason why this single material exhibits n- and p-type contact formation when the different sides of the MoSSe attach to graphene. Also, the Dirac cone of the graphene in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al. (2010)
We are also interested in metal electrode cases such as Al and Cu. For Al (111) face attachment to the S and Se faces of monolayer MoSSe [Figs. 1(f) and 1(i)], the MoSSe band intersects the Fermi level, indicating that Ohmic contact occurs between the metal and MoSSe states. The intersection can extend to the MoSSe conduction band when the S face attaches to Al, behaving as an n-type contact; p-type contact can be observed in the case of Se-face attachment, since the crossover point is mainly attributable to the valence band. Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111). In contrast to the 2L-MoSSe/graphene case, these metal contact projected band structures demonstrate the chemisorption characteristics due to significant hybridization in the band diagram. This chemisorption-like property imparts an the advantage with regard to tunneling barrier reduction at the contact region, thus improving device performance.Gong et al. (2010); Zhao and Schwingenschlögl (2022)
DFT calculations are also used to obtain the device simulation parameters and the electronic parameters MoSSe, such as bandgap, affinity, dielectric constant, and effective mass. Figure 2(a) shows the band diagram of the AAA MoSSe stacking with a relaxed lattice constant of 3.25 Å and an interlayer distance of 3.19 Å. All three layers exhibit an indirect bandgap with a valley in the conduction bands and a valley in the valence bands. The projected band structure of each MoSSe layer can be observed in Fig. 2(b). The yellow, blue, and red dots represents the projections of the top, middle, and bottom MoSSe layers, respectively, and are correlated with the colored arrows in Fig. 2(a). As the layers’ band components were separated, we observed that the bands were shifted by 0.43 eV as each MoSSe layer was added. This effect originates from the intrinsic dipole in the out-of-plane direction of MoSSe. Similar to the case of gallium nitride, we can represent this dipole effect by a spontaneous polarization term in the device simulations, which is illustrated by the two diagonal black arrows in Fig. 2(b). The term for MoSSe can be written as the product of the out-of-plane dielectric constant and the intrinsic electric field :
It is easy to obtain from the thickness of the MoSSe and the value of the out-of-plane dipole. However, a problem arises when attempting to obtain in DFT calculations, because the vacuum layer along the axis will also be included. The principle of the equivalent capacitance methodLaturia, de Put, and Vandenberghe (2018) is adopted here to separate the dielectric constants of MoSSe from those of the whole system. Carrier mobility is obtained using the Monte Carlo procedure described by Pai and Wu.Pai and Wu (2022) We design our device using MoSSe of 5 nm channel thickness to represent MoSSe multilayers, since this thickness ensures there is sufficient charge flowing in the channel region. All the parameters used in the 2D-DDCC simulation are given in Table 1. In our device simulation setting, the metal work function is set at the mid-gap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst case metal selection scenario.
| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO and AlO parameters are given inherently in 2D-DDCC.Wu |
| 1.212 eV | Average bandgap | |
| 3.97 eV | Electron affinity | |
| 667 cm/(Vs) | Electron mobility | |
| 624 cm/(Vs) | Hole mobility | |
| Electron -valley effective mass | ||
| Electron -valley effective mass | ||
| Hole -valley effective mass | ||
| Hole -valley effective mass | ||
| In-plane dielectric constant | ||
| Out-of-plane dielectric constant | ||
| cm | Spontaneous polarization | |
| SiO dielectric constant | ||
| AlO dielectric constant | ||
| 5.8 eV | HfO bandgapWu, Alivov, and Negi (2008) | |
| HfO dielectric constantKumar, Kaushik, and Negi (2014) |
Figure 3(a) is a schematic of double-gate CMOSFET device design exploiting the strong properties of MoSSe. As illustrated to the left of the black dashed line, to create a p-MOSFET, two metal electrodes are attached to the Se face to produce p-type source and drain regions, while the remaining MoSSe region between them acts like a channel area after the oxide and gate metal have been placed in the vertical direction. An n-MOSFET can be created similarly, as illustrated to the right of the dashed line, with the only difference being that in this case it is the S face, rather than the Se face, to which metals are attached to form n-type regions at source and drain. This horizontal CMOSFET design has two advantages: (1) the Janus MoSSe reduces the complexity of material selection, since only one channel material is used to form n- and p-MOSFETs; (2) this structure enables layer-by-layer build-up during the fabrication process, which makes the process more reliable.
To investigate the band and carrier characteristics of the designed MoSSe MOSFETs, the 2D-DDCC solver is used out to simulate device characteristics. Cross-sectional views of the band diagram are shown in Figs. 3(b) and 3(c) for the p- and n-MOSFETs, respectively. The strong in the MoSSe vertical direction acts like partial positive and negative charges attracting electrons and holes, causing the valence band to be pulled upward [Fig. 3(b)] and the conduction band to be bent downward [Fig. 3(c)]. Thus, the majority carrier concentration (dashed curve) surges near the interface with the metal for contacts of both polarities. Both electron and hole carrier concentrations can reach to near the order of cm when layered MoSSe serves as the channel layer. As with doping, the high carrier concentration of the metal–SeMoS contacts means that these can be regarded as p-type Ohmic contacts. The metal–SMoSe contacts can be regarded as n-type Ohmic contacts, which is in good agreement with the aforementioned DFT calculations. The difference in physical behavior between the different sides of MoSSe attachment corresponds simply to the difference in sign of here. It is also worth considering how the value of affects this phenomenon. Figure 3(d), taking an n-type MOSFET as an example, shows the effects of different values of on the band structure and carrier density at the interface. This can provide a reference for future studies of materials with different JTMD dipole strengths. As the dipole field of the material becomes stronger, a larger is generated, leading to more pronounced JTMD-induced band bending. Consequently, the accumulation of carriers becomes more prominent. Conversely, smaller values result in fewer carriers gathering at the interface. If the goal in device design is to increase carrier concentration at the contact, then selecting materials with higher values will be appropriate.
The – characteristics of several MoSSe MOSFET designs are now investigated. Figure 4(a) shows a single-gate design with gate length m, which is a common dimension adopted in experimental studies of FETs based on 2D materialsXiong et al. (2021); 2D_FET_Bench. Device modifications such as the use of double gates [Fig. 4(b)] and smaller dimension [Fig. 4(c)] are also simulated. The simulations for all devices assume an HfO gate oxide of thickness nm. Figure 4(d) shows the transfer characteristics (–) of the n- and p-MOSFETs from Figs. 4(a)–4(c). The curves with squares, circles, and diamonds correspond to the n-MOSFETs in (a), (b), and (c), respectively, and those with triangles, stars, and pentagons to the p-MOSFETs in (a), (b), and (c), respectively. The key distinction lies in the opposite orientation of in the JTMD layer. The typical – curves for the n- and p-MOSFETs indicate that all three devices exhibit satisfactory enhancement-mode MOSFET behavior. Among the three devices, the single-gate design [Figure 4(a)] has the most significant sub-threshold swing (SS) and the lowest on-current . When a bottom gate is added [Figure 4(b)], the SS drops to around 60.5 mV/dec, while increases by an order of magnitude, indicating an improvement in gate control. Having shown that both designs are feasible at a larger scale, we reduce the horizontal scale of the double-gate design to nm [Figure 4(c)] to examine the device performance at a smaller scale. The consequent increases in SS from 60.45 to 60.53 for the n-MOSFET and from 60.55 to 60.71 for the p-MOSFET demonstrate that reducing weakens the control of the gate. In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the off-state current is decreased by about an order, indicating that the loss of SS is compensated by an improvement in . All of these results show that the SS and on/off-current characteristics of the designs considered are those of a typical MOSFET device. The lower SS and higher consistently observed for all the n-MOSFETs compared with the corresponding p-MOSFETs may originate from the slightly higher mobility of electrons compared with holes.
For successful operation of a CMOSFET, an enhancement-mode MOSFET device is essential. Since the gate oxide thickness plays a crucial role in determining the operation mode of a MOSFET, a variety of oxides and different values of , are tested to determine the optimal design of the device in Fig. 5(a) in terms of the equivalent oxide thickness (EOT), and the results are shown in Figs. 5(b) and 5(c) for the n- and p-MOSFETs, respectively. Here, we define on/off state switching when . First, a SiO oxide gate of thickness 1 nm is tested, and the – result is shown by the curve with red circles. The device is clearly in the on-state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device. To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an AlO gate with nm (EOT 0.42 nm, blue triangles). However, the current remains of the order of A/cm at V, and both polarities remain in the depletion mode. Therefore, a high- material, namely, HfO, is selected to help decrease the EOT of the MOSFETs. Using HfO with nm (EOT 0.16 nm, green stars) gives a superior enhancement-mode device for both n- and p-MOSFETs. Finally, we optimize the of the HfO to 2 nm (EOT 0.31 nm, purple diamonds) to prevent current leakage through the thin gate oxide. As can be seen from Figs. 5(b) and 5(c), both of the – curves for this optimized value of intersect the off-state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased. These results for HfO with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm. For this optimal MOSFET design based on HfO with nm, the SS as extracted from Figs. 5(b) and 5(c) is about 64 mV/dec. Since such low- devices may encounter several processing difficulties, simulation results for a more easily fabricated thick HfO gate with nm (EOT 1.6 nm, black squares) are also given in Figs. 5(b) and 5(c). Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 1.2 V and 1.4 V for n- and p-MOSFETs, respectively. All in all, the – plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
To delve further into the on/off characteristics of the MOSFET devices, we plot the majority carrier concentration of the optimal p- and n-MOSFET design (HfO gate with nm) in Fig. 6. The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm (green color). The carrier concentration gradually decays as the position moves away from the contact region. Owing to the 10 nm gate metal isolation design, two bump-like regions of high carrier concentration appear next to the source and drain areas. The thicker oxide eliminates the dipole effect of MoSSe, and thus carrier gas remains. Taking the p-MOSFET (top two plots) as an example, we can see that in with the application of a negative 0.5 V gate bias, holes are attracted into the channel region, forming a strong inversion channel layer and causing the device to turn on. By contrast, when a positive 0.5 V gate bias is applied, holes are depleted, and the inversion layer disappears, indicating an off-state device. A similar result is also obtained for the n-MOSFET, with the only difference being that a strong inversion layer appears when V and the inversion layer vanishes when a V bias is applied.
IV Conclusions
We have introduced a novel CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe. Our investigation has encompassed diverse heterostructures, including bilayer MoSSe/graphene and MoSSe/Al (111) face and MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT. We have also examined the interactions between MoSSe, graphene, and Al and Cu electrodes, identifying characteristics such as Schottky barriers and Ohmic contacts. Our design leverages the strong dipole field of MoSSe to create n- and p-type contact regions, thereby forming a CMOSFET with simplified material selection. The proposed horizontal CMOSFET architecture offers advantages in fabrication reliability and integration. Device simulations have provided insights into the spontaneous polarization caused by band bending and the related carrier distribution, with an optimized EOT 0.312 nm giving a favorable threshold between the depletion and enhancement modes of the MOSFET device. Majority carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits. On the basis of the surface polarization strategy adopted here, it might be possible to form a MOSFET from any JTMD, provided that the 2D material possesses a strong intrinsic out-of-plane dipole. Regardless, this work has introduced a promising approach to CMOSFET design, offering valuable insights into materials, interfaces, and device characteristics for future electronic applications.
Acknowledgements.
This work was supported by the National Science and Technology Council under Grant Nos. NSTC 112-2221-E-002-215-MY3, NSTC 112-2923-E-002-002, and 12-2221-E-002-214-MY3.AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose
Author Contributions
Yun-Pin Chiu: Conceptualization (equal); Data curation (lead);Formal analysis (equal); Methodology (equal); Validation (lead); Visualization(lead); Writing – original draft (lead). Hsin-Wen Huang: Conceptualization (equal); Data curation (support); Formal analysis (equal); Methodology (equal); Validation (support). Yuh-Renn Wu: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Software(lead); Supervision (lead); Writing – review and editing (lead).
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@
 Janus transition metal dichalcogenides (JTMDs) have attracted much attention because of their outstanding electronic and optical properties.
-The additional out-of-plane dipole in JTMDs can form n- and p-like Ohmic contacts, and this may be used in device applications such as pin diodes and photovoltaic cells.
+The additional out-of-plane dipole in JTMDs can form n- and p-like contacts, and this may be used in device applications such as pin diodes and photovoltaic cells.
 In this study, we exploit this property to design n- and p-type metal–oxide–semiconductor field effect transistors (MOSFETs).
 First, we use density-functional theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe.
-The intrinsic dipole of MoSSe causes band bending at both the metal/MoSSe and MoSSe/metal interfaces, resulting in electron and hole accumulation to form n- and p-type Ohmic contact regions.
-We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, on the basis of which we design complementary MOSFETs.
+The intrinsic dipole of MoSSe causes band bending at both the metal/SMoSe and SMoSe/metal interfaces, resulting in electron and hole accumulation to form p-type, n-type, and even Ohmic contact regions.
+We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, based on which we design complementary MOSFETs.
 Our results demonstrate that JTMDs can be used to make n- and p-MOSFETs in the same layer without the need for any extra doping.
@@
 (2014) These properties make 2D TMDs suitable for various device applicationsLiu et al.
-(2021); Cho, Park, and Kwon (2015); Ying-Chuan_KMC; Xiong et al.
-(2021); 2D_Photo, especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs).
+(2021); Cho, Park, and Kwon (2015); Chen et al.
+(2023); Xiong et al.
+(2021); Zhao et al.
+(2021), especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs).
 Liu et al.
-(2021); Chhowalla, Jena, and Zhang (2016b); 2D_FET_Pin-Fang These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials.
-For example, MoS and WS are two commonly selected materials for n-type channels, and -MoTe and WSe are both good candidates for p-type channels.Radisavljevic et al.
+(2021); Chhowalla, Jena, and Zhang (2016b); Chen, Chen, and Wu (2022) These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials.
+For example, MoS2 and WS2 are two commonly selected materials for n-type channels, and -MoTe2 and WSe2 are both good candidates for p-type channels.Radisavljevic et al.
 (2011); Pezeshki et al.
@@
 (2019); Zheng et al.
-(2021) For an FET with SB contacts, carriers must overcome the energy barrier to transport from metal into the TMD channel region, impairing the device function by lowering the injection efficiency.Tung (2014) A typical technique used to reduce contact resistance in modern semiconductor technology is substitutional doping.Younas, Zhou, and Hinkle (2023) However, there are better ways to apply such a doping method to 2D materials, since their characteristics will be heavily altered after doping.
+(2021) For a FET with SB contacts, carriers must overcome the energy barrier to transport from metal into the TMD channel region, impairing the device function by lowering the injection efficiency.Tung (2014) A typical technique used to reduce contact resistance in modern semiconductor technology is substitutional doping.Younas, Zhou, and Hinkle (2023) However, there are better ways to apply such a doping method to 2D materials since their characteristics will be heavily altered after doping.
 Additionally, the absence of dangling bonds on the surface of a 2D material makes solid interface connections with metal challenging, resulting in larger contact resistance.Allain et al.
@@
 For n-FETs, a material with a smaller work function is needed, whereas a metal with a high work function is often selected for p-FETs.
-Several studies have shown that TMD/metal heterostructures such as MoS/Cu, WS/Cu, and WSe/Ti could provide n-type SB contacts,Kang et al.
-(2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe/Pd and WSe/Au.Kang et al.
+Several studies have shown that TMD/metal heterostructures such as MoS2/Cu, WS2/Cu, and WSe2/Ti could provide n-type SB contacts,Kang et al.
+(2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe2/Pd and WSe2/Au.Kang et al.
 (2014); Kong et al.
@@
 (2021) Furthermore, two contact metals need to be considered separately, resulting in higher processing complexity and a more time-consuming fabrication procedure.
-Recently, ambipolar TMDs such as MoTeLin et al.
-(2014) and ReSeLee et al.
+Recently, ambipolar TMDs such as MoTe2Lin et al.
+(2014) and ReSe2Lee et al.
 (2019) have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
@@
 (2022, 2020); Kong et al.
-(2023) Also, simulation studies have shown that this intrinsic dipole feature of JTMDs can enable the formation of both n- and p-type contacts simultaneously without the need for additional doping when different sides are attached to graphene,Chen et al.
+(2023) Also, simulation studies have shown that this intrinsic dipole feature of JTMDs can enable the formation of both n- and p-type contacts simultaneously without the need for additional doping when different sides are attached to graphene (Gr),Chen et al.
 (2022b); Palsgaard et al.
@@
 In this study, we propose a CMOSFET design concept utilizing the intrinsic dipole of a JTMD.
-The material MoSSe is selected for investigation, since it was the first JTMD and is the most commonly synthesized.Lu et al.
+The material MoSSe was selected for investigation since it was the first JTMD and is the most commonly synthesized.Lu et al.
 (2017); Trivedi et al.
 (2020) The remainder of this paper is organized as follows.
-First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, bilayer MoSSe (2L-MoSSe)/graphene and MoSSe/Al (111) face and MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT).
+First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, Gr/trilayer AA′A stacking MoSSe/Gr and A′A stacking bilayer MoSSe(2L-MoSSe)/Al (111) face and 2L-MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT).
 Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson–drift–diffusion solver (2D-DDCC) developed in our laboratory.Wu
@@
 (2010) An Monkhorst–Pack sampling is used.
-The geometry is refined until the maximal residual forces are less than 0.015 eV/Å, at which point MoSSe heterostructures are subjected to the dipole correction to negate the artificial electric field caused by their asymmetric structures.
+The geometry is refined until the maximal residual forces are less than 0.015 eV/Å.
+At this point, MoSSe heterostructures are subjected to dipole correction to negate the artificial electric field caused by their asymmetric structures.
 The vacuum region is set to 20 Å to prevent interactions between nearby supercells.
-For MoSSe contact simulation, we attach a 1 supercell of graphene to a two-layer AA stacking supercell of MoSSe, A supercell of six layers of (111) Al to a 1 supercell of MoSSe, and a 1 supercell of six layers of (111) Cu to a supercell of MoSSe.
-A simulation model is created for each metal’s energetically advantageous surface,Skriver and Rosengaard (1992) and four outward metal layers are fixed in the relaxation process.
-For layered MoSSe, we construct an AAA stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here.
+For MoSSe contact simulation, we attach a 1 supercell of Gr to a 2L-MoSSe supercell.
+A supercell of six layers of (111) Al to a 1 2L-MoSSe supercell, and a 1 supercell of six layers of (111) Cu to a 2L-MoSSe supercell.
+A simulation model is created for each metal’s energetically advantageous surface,Skriver and Rosengaard (1992), and four outward metal layers are fixed in the relaxation process.
+For layered MoSSe, we construct an AA′A stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here.
 The only difference is that the mesh is set to in the MoSSe band structure calculation.
 II.2 Device simulation
-The 2D-DDCC solver is used to examine the device performance.
+The 2D-DDCC solver is used to examine the device’s performance.
 This solver was developed in our laboratory and has been used on many different kinds of devices.
@@
 where is the dielectric constant, is the electric potential, is the acceptor’s doping concentration, is the donor’s doping concentration, is the electron concentration, is the hole concentration, and are the current densities of electrons and holes, respectively, and and are the electron and hole diffusion coefficients, respectively.
-This solver serves as the basis for all of the device simulations discussed in this article, with the exception of the MoSSe channel region calculations.
-III Results and discussion
-To study the interactions between MoSSe and contacts, we construct the S and Se sides of the MoSSe heterostructure with graphene and two common electrode metals (namely, Al and Cu).
-Figure 1(a) shows the calculated heterostructure systems that can form low SB or Ohmic contacts.
-The top views are shown in Figs.
-1(b) and 1(c), where the lattice mismatches are 1.11%, 0.44%, and 2.97% for the 2L-MoSSe/graphene, MoSSe/Al, and MoSSe/Cu systems, respectively.
-Figures 1(e)–1(j) show the projected band structures, and Figs.
-1(e)–1(g) representing the S-face contact of MoSSe with the electrode material, and Figs.
-1(h)–1(j) representing the Se-face attachment.
-As can be observed, if graphene contacts a two-layer MoSSe material on the Se side [Fig.
-1(e)], then the Fermi level of the entire system is 0.1 eV lower than the observable conduction band of the MoSSe material, suggesting an n-type Schottky barrier.
-On the other hand, if graphene contacts a two-layer MoSSe material on the S side [Fig.
-1(h)], then we see a p-type Ohmic contact.
-These results are consistent with the results of previous JTMD contact studies.Palsgaard et al.
-(2018); Chen et al.
-(2022b) According to our DFT calculation, monolayer MoSSe has a strong V/cm intrinsic dipole field in the out-of-plane direction, which may be the reason why this single material exhibits n- and p-type contact formation when the different sides of the MoSSe attach to graphene.
-Also, the Dirac cone of the graphene in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al.
-(2010)
-We are also interested in metal electrode cases such as Al and Cu.
-For Al (111) face attachment to the S and Se faces of monolayer MoSSe [Figs.
-1(f) and 1(i)], the MoSSe band intersects the Fermi level, indicating that Ohmic contact occurs between the metal and MoSSe states.
-The intersection can extend to the MoSSe conduction band when the S face attaches to Al, behaving as an n-type contact; p-type contact can be observed in the case of Se-face attachment, since the crossover point is mainly attributable to the valence band.
-Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111).
-In contrast to the 2L-MoSSe/graphene case, these metal contact projected band structures demonstrate the chemisorption characteristics due to significant hybridization in the band diagram.
-This chemisorption-like property imparts an the advantage with regard to tunneling barrier reduction at the contact region, thus improving device performance.Gong et al.
-(2010); Zhao and Schwingenschlögl (2022)
-DFT calculations are also used to obtain the device simulation parameters and the electronic parameters MoSSe, such as bandgap, affinity, dielectric constant, and effective mass.
-Figure 2(a) shows the band diagram of the AAA MoSSe stacking with a relaxed lattice constant of 3.25 Å and an interlayer distance of 3.19 Å.
-All three layers exhibit an indirect bandgap with a valley in the conduction bands and a valley in the valence bands.
-The projected band structure of each MoSSe layer can be observed in Fig.
-2(b).
-The yellow, blue, and red dots represents the projections of the top, middle, and bottom MoSSe layers, respectively, and are correlated with the colored arrows in Fig.
-2(a).
-As the layers’ band components were separated, we observed that the bands were shifted by 0.43 eV as each MoSSe layer was added.
-This effect originates from the intrinsic dipole in the out-of-plane direction of MoSSe.
-Similar to the case of gallium nitride, we can represent this dipole effect by a spontaneous polarization term in the device simulations, which is illustrated by the two diagonal black arrows in Fig.
-2(b).
-The term for MoSSe can be written as the product of the out-of-plane dielectric constant and the intrinsic electric field :
-It is easy to obtain from the thickness of the MoSSe and the value of the out-of-plane dipole.
-However, a problem arises when attempting to obtain in DFT calculations, because the vacuum layer along the axis will also be included.
-The principle of the equivalent capacitance methodLaturia, de Put, and Vandenberghe (2018) is adopted here to separate the dielectric constants of MoSSe from those of the whole system.
-Carrier mobility is obtained using the Monte Carlo procedure described by Pai and Wu.Pai and Wu (2022) We design our device using MoSSe of 5 nm channel thickness to represent MoSSe multilayers, since this thickness ensures there is sufficient charge flowing in the channel region.
-All the parameters used in the 2D-DDCC simulation are given in Table 1.
-In our device simulation setting, the metal work function is set at the mid-gap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst case metal selection scenario.
-| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO and AlO parameters are given inherently in 2D-DDCC.Wu |
+This solver serves as the basis for all of the device simulations discussed in this article, except the MoSSe channel region calculations.
+| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO2 and Al2O3 parameters are given inherently in 2D-DDCC.Wu |
 | 1.212 eV | Average bandgap | |
 | 3.97 eV | Electron affinity | |
-| 667 cm/(Vs) | Electron mobility | |
-| 624 cm/(Vs) | Hole mobility | |
+| 667 cm2/(Vs) | Electron mobility | |
+| 624 cm2/(Vs) | Hole mobility | |
 | Electron -valley effective mass | ||
@@
 | Out-of-plane dielectric constant | ||
-| cm | Spontaneous polarization | |
-| SiO dielectric constant | ||
-| AlO dielectric constant | ||
-| 5.8 eV | HfO bandgapWu, Alivov, and Negi (2008) | |
-| HfO dielectric constantKumar, Kaushik, and Negi (2014) |
+| cm-2 | Spontaneous polarization | |
+| SiO2 dielectric constant | ||
+| Al2O3 dielectric constant | ||
+| 5.8 eV | HfO2 bandgapWu, Alivov, and Negi (2008) | |
+| HfO2 dielectric constantKumar, Kaushik, and Negi (2014) |
+III Results and discussion
+DFT calculations are also used to obtain the device simulation parameters and the electronic parameters MoSSe, such as bandgap, affinity, dielectric constant, and effective mass.
+Figure 1(a) shows the band diagram of the AA′A MoSSe stacking with a relaxed lattice constant of 3.25 Å and an interlayer distance of 3.19 Å.
+All three layers exhibit an indirect bandgap with a valley in the conduction bands and a valley in the valence bands.
+The projected band structure of each MoSSe layer can be observed in Fig.
+1(b).
+The yellow, blue, and red dots represent the projections of the top, middle, and bottom MoSSe layers, respectively, and are correlated with the colored arrows in Fig.
+1(a).
+As the layers’ band components were separated, we observed that the bands were shifted by 0.43 eV as each MoSSe layer was added.
+This effect originates from the intrinsic dipole in the out-of-plane direction of MoSSe.
+Similar to the case of gallium nitride, we can represent this dipole effect by a spontaneous polarization term in the device simulations, which is illustrated by the two diagonal black arrows in Fig.
+1(b).
+The term for MoSSe can be written as the product of the out-of-plane dielectric constant and the intrinsic electric field :
+It is easy to obtain from the thickness of the MoSSe and the value of the out-of-plane dipole.
+However, a problem arises when attempting to obtain in DFT calculations, because the vacuum layer along the axis will also be included.
+The principle of the equivalent capacitance methodLaturia, de Put, and Vandenberghe (2018) is adopted here to separate the dielectric constants of MoSSe from those of the whole system.
+Carrier mobility is obtained using the Monte Carlo procedure described by Pai and Wu.Pai and Wu (2022)
+The aforementioned also influences the interactions between MoSSe and contacts.
+Initially, we construct the Gr/AA′A-MoSSe/Gr heterostructure, with the S and Se sides attached to the Gr electrodes, resulting in a lattice mismatch of 1.11% between MoSSe and Gr.
+In Figure 2(b), the projected band structures of the Gr/AA′A-MoSSe/Gr system are illustrated.
+Notably, when Gr contacts a three-layer MoSSe material on the Se side (depicted in red), the Fermi level of the entire system is observed to be 0.38 eV lower than the conduction band of the MoSSe material, indicating an n-type Schottky barrier.
+Conversely, when Gr contacts a three-layer MoSSe material on the S side (depicted in green), a p-type Ohmic contact is observed at the Gamma point.
+These findings align with previous studies on JTMD contacts.Palsgaard et al.
+(2018); Chen et al.
+(2022b) According to our DFT calculation, monolayer MoSSe exhibits a strong V/cm intrinsic dipole field in the out-of-plane direction.
+This characteristic may explain why this single material exhibits both n- and p-type contact formations when different sides of MoSSe attach to Gr.
+Furthermore, in Figure 2(c), the effect of the intrinsic dipole on the system is depicted; the pulls the Gr downward as the Se side contacts the Gr.
+Consequently, the Fermi level becomes higher than the Dirac cone, resulting in intrinsically n-doped Gr.
+Due to the tilting band edge caused by , the energy difference between the conduction band edge and the Fermi level is smaller than the valence band edge offset at the Se/Gr side, suggesting the formation of an n-region.
+On the other hand, the S-side attachment elevates the band energy of the Gr, with the valence band edge intersecting the Fermi level, indicating an intrinsic p-doped Gr and a p-contact region.
+Nevertheless, the Dirac cone of both Gr in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al.
+(2010)
+We also simulate metal contact systems due to the interest in metal electrode cases, where common electrode materials such as Al and Cu are selected.Allain et al.
+(2015); Kang et al.
+(2014); Ouyang, Xiong, and Jing (2018); Tang et al.
+(2019); Zheng et al.
+(2021) Fig.
+3(a) illustrates the simulated 2L-MoSSe/Metal heterostructure and the top views are shown in Figs.
+3(b) and 3(c), where the lattice mismatches are 0.44%, and 2.97% for the 2L-MoSSe/Al and 2L-MoSSe/Cu systems.
+Figures 3(d)–3(g) show the projected band structures, and Figs.
+3(d)–3(f) representing the Se-face contact of MoSSe with the electrode material, and Figs.
+3(e)–3(g) representing the S-face attachment.
+For Al (111) face attachment to the Se and S faces of bilayer MoSSe (2L-MoSSe) [Figs.
+3(d) and 3(e)], the 2L-MoSSe band intersects the Fermi level, indicating that Ohmic contact occurs between the metal and 2L-MoSSe states.
+In contrast to the 2L-MoSSe/Gr cases, these metal contact projected band structures demonstrate the chemisorption characteristics due to significant hybridization in the band diagram.
+Different from typical TMDs, the internal electric field in JTMD plays a crucial role in augmenting the density of interface states.
+Specifically, in MoSSe, this electric field fosters the generation of electronic energy levels at the metal-semiconductor interface via chemisorption-like bonding, leading to an increased density of interface states between the two materials.
+Zhao and Schwingenschlögl (2022) These additional bonding states, resulting from in MoSSe, are delineated in density of states (DOS) plots for all systems, indicated by a dotted frame.Gong et al.
+(2010) Due to the additional bonding states caused by chemical adsorption, it becomes challenging to correlate the band edge of MoSSe closest to the metal layer.
+We correlate the band edge distribution of the closest-to-metal layer of MoSSe (1L MoSSe) based on the direction of previously mentioned and the peak present at the Ev edge of MoSSe layers, and depict the corresponding band edge with DOS on the right side of Figs.
+3(d) and (e).
+When the Se face contacts with Al [Fig.
+3(d)], it can be observed that the Fermi level coincides with the outer MoSSe (2L MoSSe) at the valence band edge.
+Due to the upward direction of , the energy band of the inner layer of MoSSe (1L MoSSe) shifts towards lower energy.
+Supported by the peak at the Ev edge, it can be inferred that the position of the Fermi level is closer to the conduction band, leaning towards an n-type contact.
+Conversely, when the S face contacts with Al [Fig.
+3(e)], with the downward direction of , the energy band of the 1L MoSSe shifts towards higher energy.
+A comparison reveals that the Fermi level is situated in the middle of the band, leaning towards a p-type contact.
+Although we correlate the n- or p-type contact in both systems, there are numerous bonding states present at the gap position of 1L MoSSe, indicating the Ohmic contact formation.
+In other words, both electrons and holes can freely flow through this low-tunneling barrier Ohmic contact, thereby enhancing device performance.Gong et al.
+(2010); Zhao and Schwingenschlögl (2022) Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111) [Fig.
+3(f) and (g)], indicating Cu can also bring Ohmic contact properties as 2L-MoSSe attches.
+The chemisorption-like property imparts the advantage concerning tunneling barrier reduction at the contact region, thus improving device performance.
+Zhao and Schwingenschlögl (2022); Gong et al.
+(2010)
 Figure 4(a) is a schematic of double-gate CMOSFET device design exploiting the strong properties of MoSSe.
-As illustrated to the left of the black dashed line, to create a p-MOSFET, two metal electrodes are attached to the Se face to produce p-type source and drain regions, while the remaining MoSSe region between them acts like a channel area after the oxide and gate metal have been placed in the vertical direction.
-An n-MOSFET can be created similarly, as illustrated to the right of the dashed line, with the only difference being that in this case it is the S face, rather than the Se face, to which metals are attached to form n-type regions at source and drain.
-This horizontal CMOSFET design has two advantages: (1) the Janus MoSSe reduces the complexity of material selection, since only one channel material is used to form n- and p-MOSFETs; (2) this structure enables layer-by-layer build-up during the fabrication process, which makes the process more reliable.
-To investigate the band and carrier characteristics of the designed MoSSe MOSFETs, the 2D-DDCC solver is used out to simulate device characteristics.
+As illustrated to the left of the black dashed line, to create a p-MOSFET, two metal electrodes are attached to the S face to produce a p-type source and drain regions, while the remaining MoSSe region between them acts like a channel area after the oxide and gate metal have been placed in the vertical direction.
+An n-MOSFET can be created similarly, as illustrated to the right of the dashed line, with the only difference being that in this case, it is the Se face, rather than the S face, to which metals are attached from n-type regions at the source and drain.
+This horizontal CMOSFET design has two advantages: (1) the Janus MoSSe reduces the complexity of material selection since only one channel material is used to form n- and p-MOSFETs; (2) this structure enables layer-by-layer build-up during the fabrication process, which makes the process more reliable.
+To investigate the electron band characteristics of the designed MoSSe MOSFETs, the 2D-DDCC solver is used to simulate device characteristics.
+We design our device using MoSSe of 5 nm channel thickness to represent MoSSe multilayers since this thickness ensures there is sufficient charge flowing in the channel region.
+All the parameters used in the 2D-DDCC simulation are given in Table 1.
+In our device simulation setting, the metal work function is set at the mid-gap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst-case metal selection scenario.
 Cross-sectional views of the band diagram are shown in Figs.
-3(b) and 3(c) for the p- and n-MOSFETs, respectively.
-The strong in the MoSSe vertical direction acts like partial positive and negative charges attracting electrons and holes, causing the valence band to be pulled upward [Fig.
-3(b)] and the conduction band to be bent downward [Fig.
-3(c)].
-Thus, the majority carrier concentration (dashed curve) surges near the interface with the metal for contacts of both polarities.
-Both electron and hole carrier concentrations can reach to near the order of cm when layered MoSSe serves as the channel layer.
-As with doping, the high carrier concentration of the metal–SeMoS contacts means that these can be regarded as p-type Ohmic contacts.
-The metal–SMoSe contacts can be regarded as n-type Ohmic contacts, which is in good agreement with the aforementioned DFT calculations.
+4(b) and 4(c) for the p- and n-MOSFETs, respectively.
+The strong in the MoSSe vertical direction acts like partial positive and negative charges attracting electrons and holes.
+The negative polarization charge causes the Ec and Ev to be pulled upward [Fig.
+4(b)] and negative polarization charge casues the Ec and Ev band to be bent downward [Fig.
+4(c)] at the interface of the MoSSe region.
+This band edge bending, originating from , plays a pivotal role in facilitating the formation of ohmic contacts within the device, which is in good agreement with the aforementioned DFT calculations.
 The difference in physical behavior between the different sides of MoSSe attachment corresponds simply to the difference in sign of here.
-It is also worth considering how the value of affects this phenomenon.
-Figure 3(d), taking an n-type MOSFET as an example, shows the effects of different values of on the band structure and carrier density at the interface.
-This can provide a reference for future studies of materials with different JTMD dipole strengths.
-As the dipole field of the material becomes stronger, a larger is generated, leading to more pronounced JTMD-induced band bending.
-Consequently, the accumulation of carriers becomes more prominent.
-Conversely, smaller values result in fewer carriers gathering at the interface.
-If the goal in device design is to increase carrier concentration at the contact, then selecting materials with higher values will be appropriate.
+In contrast, when simulating TMDs lacking this intrinsic dipole, where is set to zero (dot lines labeled "w/o "), the same device structure is employed.
+In the absence of assistance, the formation of Ohmic contacts falls short, resulting in the Schottky barrier at the interface.
 The – characteristics of several MoSSe MOSFET designs are now investigated.
 Figure 5(a) shows a single-gate design with gate length m, which is a common dimension adopted in experimental studies of FETs based on 2D materialsXiong et al.
-(2021); 2D_FET_Bench.
+(2021); Sebastian et al.
+(2021).
 Device modifications such as the use of double gates [Fig.
-4(b)] and smaller dimension [Fig.
-4(c)] are also simulated.
-The simulations for all devices assume an HfO gate oxide of thickness nm.
+5(b)] and smaller dimension [Fig.
+5(c)] are also simulated.
+The simulations for all devices assume an HfO2 gate oxide of thickness nm.
 Figure 5(d) shows the transfer characteristics (–) of the n- and p-MOSFETs from Figs.
-4(a)–4(c).
+5(a)–5(c).
 The curves with squares, circles, and diamonds correspond to the n-MOSFETs in (a), (b), and (c), respectively, and those with triangles, stars, and pentagons to the p-MOSFETs in (a), (b), and (c), respectively.
@@
 Having shown that both designs are feasible at a larger scale, we reduce the horizontal scale of the double-gate design to nm [Figure 5(c)] to examine the device performance at a smaller scale.
-The consequent increases in SS from 60.45 to 60.53 for the n-MOSFET and from 60.55 to 60.71 for the p-MOSFET demonstrate that reducing weakens the control of the gate.
-In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the off-state current is decreased by about an order, indicating that the loss of SS is compensated by an improvement in .
+The proximity of the SS values in the n-MOSFET suggests that reducing to 200 nm would not significantly impact the control of the gate in this design.
+In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the off-state current is decreased by about an order, indicating an improvement in .
 All of these results show that the SS and on/off-current characteristics of the designs considered are those of a typical MOSFET device.
 The lower SS and higher consistently observed for all the n-MOSFETs compared with the corresponding p-MOSFETs may originate from the slightly higher mobility of electrons compared with holes.
-For successful operation of a CMOSFET, an enhancement-mode MOSFET device is essential.
+For the successful operation of a CMOSFET, an enhancement-mode MOSFET device is essential.
 Since the gate oxide thickness plays a crucial role in determining the operation mode of a MOSFET, a variety of oxides and different values of , are tested to determine the optimal design of the device in Fig.
-5(a) in terms of the equivalent oxide thickness (EOT), and the results are shown in Figs.
-5(b) and 5(c) for the n- and p-MOSFETs, respectively.
+6(a) in terms of the equivalent oxide thickness (EOT), and the results are shown in Figs.
+6(b) and 6(c) for the n- and p-MOSFETs, respectively.
 Here, we define on/off state switching when .
-First, a SiO oxide gate of thickness 1 nm is tested, and the – result is shown by the curve with red circles.
+First, a SiO2 oxide gate of thickness 1 nm is tested, and the – result is shown by the curve with red circles.
 The device is clearly in the on-state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device.
-To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an AlO gate with nm (EOT 0.42 nm, blue triangles).
+To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an Al2O3 gate with nm (EOT 0.42 nm, blue triangles).
 However, the current remains of the order of A/cm at V, and both polarities remain in the depletion mode.
-Therefore, a high- material, namely, HfO, is selected to help decrease the EOT of the MOSFETs.
-Using HfO with nm (EOT 0.16 nm, green stars) gives a superior enhancement-mode device for both n- and p-MOSFETs.
-Finally, we optimize the of the HfO to 2 nm (EOT 0.31 nm, purple diamonds) to prevent current leakage through the thin gate oxide.
+Therefore, a high- material, namely, HfO2, is selected to help decrease the EOT of the MOSFETs.
+Using HfO2 with nm (EOT 0.16 nm, green stars) gives a superior enhancement-mode device for both n- and p-MOSFETs.
+Finally, we optimize the of the HfO2 to 2 nm (EOT 0.31 nm, purple diamonds) to prevent current leakage through the thin gate oxide.
 As can be seen from Figs.
-5(b) and 5(c), both of the – curves for this optimized value of intersect the off-state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased.
-These results for HfO with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm.
-For this optimal MOSFET design based on HfO with nm, the SS as extracted from Figs.
-5(b) and 5(c) is about 64 mV/dec.
-Since such low- devices may encounter several processing difficulties, simulation results for a more easily fabricated thick HfO gate with nm (EOT 1.6 nm, black squares) are also given in Figs.
-5(b) and 5(c).
-Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 1.2 V and 1.4 V for n- and p-MOSFETs, respectively.
+6(b) and 6(c), both of the – curves for this optimized value of intersect the off-state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased.
+These results for HfO2 with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm.
+For this optimal MOSFET design based on HfO2 with nm, the SS as extracted from Figs.
+6(b) and 6(c) is about 63 mV/dec.
+Since such low- devices may encounter several processing difficulties, simulation results for a more easily fabricated thick HfO2 gate with nm (EOT 1.6 nm, black squares) are also given in Figs.
+6(b) and 6(c).
+Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 2.1 V and 1.8 V for n- and p-MOSFETs, respectively.
 All in all, the – plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
-To delve further into the on/off characteristics of the MOSFET devices, we plot the majority carrier concentration of the optimal p- and n-MOSFET design (HfO gate with nm) in Fig.
-6.
-The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm (green color).
+It is also worth considering how the value of affects this carrier accumulation phenomenon at the channel MoSSe region.
+To reveal the impact of the intrinsic dipole’s magnitude on device performance, we calculate the two-dimensional electron gas (2DEG) in the n-MOSFET device with a 200 nm length channel, shown in [Figure 7(a)], with various assumed values ranging from cm-2 to cm-2.
+These values can be regarded as JTMDs with different intrinsic dipole strengths, providing insights into how the magnitude influences the device’s 2DEG.
+The 2DEG corresponding to MoSSe’s is specified in the black dot frame.
+SiO2 ( and nm) and HfO2 ( and nm) were selected as representative low- and high- gate oxides for this study.
+To isolate the impact of , the calculations were performed at zero bias, corresponding to the equilibrium state.
+[Figure 7(b)] illustrates the influence of on the 2DEG concentration.
+In cases with high EOT (SiO2, nm), 2DEG exhibits a nearly linear increase with increasing .
+This linearity arises because the high EOT allows carriers to reach saturation within the investigated range ( to cm-2).
+However, as EOT decreases, the starting point of the linear region shifts towards higher values.
+For instance, the linear region begins at cm-2 for HfO2 with nm and cm-2 for SiO2 with nm.
+The value at the onset of the linear regime serves as an indicator for achieving enhanced mode devices.
+The HfO2 design with nm has the lowest EOT, resulting in a linear regime starting beyond cm-2.
+This suggests that MoSSe (marked by the brown dotted box) will operate in enhanced mode due to the initially low 2DEG concentration.
+Conversely, increasing the HfO2 thickness to 3 nm lowers the EOT and brings the linear region starting point to cm-2.
+This indicates a depletion mode behavior for MoSSe with HfO2 ( nm) due to the higher initial induced 2DEG.
+The influence of extends to the IV characteristics of MOSFETs.
+To focus on its impact on the linear 2DEG region, we examined devices with values of , , and cm-2, all designed with a 10 nm thick HfO2.
+As shown in Figure 7(c), an increased shifts the transfer IV curve towards more negative voltages.
+This shift reflects the need for a stronger gate voltage to deplete the augmented 2DEG in the channel.
+Moreover, the higher carrier concentration resulting from a higher leads to a greater on-current (Ion).
+Additionally, can influence carrier injection at the contacts.
+For strengths of and cm-2 value of , the device also has a larger gm value due to a smaller contact resistance.
+However, if falls below a certain threshold, devices display non-ideal IV curves near saturation (highlighted by the purple dotted box).
+As detailed in Figure 4(c), insufficient compromises Ohmic contacts, resulting in increased resistance from an additional barrier at the contact regions.
+This issue can be mitigated by selecting a metal with a work function closer to the band edge to offset the weaker band bending at the contact interface.
+For p-MOSFETs, the same principles apply.
+To understand the distribution of carriers in the device, we export the two-dimensional distribution map of the majority carrier of the device under different operating conditions in Figure 8, which demonstrates the on/off characteristics of the p- and n-MOSFET devices with HfO2 gate with nm.
+The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm-3 (green color).
 The carrier concentration gradually decays as the position moves away from the contact region.
@@
 The thicker oxide eliminates the dipole effect of MoSSe, and thus carrier gas remains.
-Taking the p-MOSFET (top two plots) as an example, we can see that in with the application of a negative 0.5 V gate bias, holes are attracted into the channel region, forming a strong inversion channel layer and causing the device to turn on.
-By contrast, when a positive 0.5 V gate bias is applied, holes are depleted, and the inversion layer disappears, indicating an off-state device.
-A similar result is also obtained for the n-MOSFET, with the only difference being that a strong inversion layer appears when V and the inversion layer vanishes when a V bias is applied.
+Taking the p-MOSFET as an example, we can see that with the application of a negative 0.5 V gate bias in Fig.
+8(a), holes are attracted into the channel region, forming a strong inversion channel layer and causing the device to turn on.
+By contrast, when a positive 0.5 V gate bias is applied in Fig.
+8(b), holes are depleted, and the inversion layer disappears, indicating an off-state device.
+A similar result is also obtained for the n-MOSFET, with the only difference being that a strong inversion layer appears when V (Fig.
+8(c)) and the inversion layer vanishes when a V bias is applied (Fig.
+8(d)).
+Figure 9 demonstrates the influence of the intrinsic dipole on the n-MOSFET channel region using a 2D simulation of a double-gate n-MOSFET device operating in both depletion mode (HfO2, nm) and enhancement mode (HfO2, nm).
+An additional dataset labeled "Without " is provided, representing traditional TMD materials, to highlight the impact of on carrier distribution in the MoSSe channel.
+All biases are zeroed to observe carrier distribution under equilibrium conditions.
+In configurations lacking an intrinsic dipole (as shown in Figs.
+9(a) and (c)), no 2DEG is observed across the channel region, with the electron concentration remaining around cm-3 regardless of oxide thickness.
+In contrast, when an intrinsic dipole is present in the channel region, it acts as a built-in electric field within the channel layer.
+This field, influenced by , is depicted in Figs.
+9(b) and (d).
+As a result, electrons align with the direction of , forming a 2DEG akin to a conventional GaN-based n-FET channel.
+Any JTMD with a pronounced will display this characteristic 2DEG formation.
+The intensity of the 2DEG is mainly dictated by the EOT.
+A value of EOT= 0.16 nm corresponds to the enhancement mode (as seen in Fig.
+9(b)).
+As EOT expands to 1.6 nm (illustrated in Fig.
+9(d)), the electron concentration at the channel interface experiences a sharp increase, causing a shift from enhancement mode to depletion mode.
+In the case of p-MOSFETs, the majority carrier distribution mirrors that of n-MOSFETs, yielding analogous outcomes.
 IV Conclusions
 We have introduced a novel CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe.
-Our investigation has encompassed diverse heterostructures, including bilayer MoSSe/graphene and MoSSe/Al (111) face and MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT.
-We have also examined the interactions between MoSSe, graphene, and Al and Cu electrodes, identifying characteristics such as Schottky barriers and Ohmic contacts.
+Our investigation has encompassed diverse heterostructures, including Gr/AA′A-MoSSe/Gr and 2L-MoSSe/Al (111) face and 2L-MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT.
+We have also examined the interactions between MoSSe, Gr, and Al and Cu electrodes, identifying characteristics such as Schottky barriers and Ohmic contacts.
 Our design leverages the strong dipole field of MoSSe to create n- and p-type contact regions, thereby forming a CMOSFET with simplified material selection.
@@
 Device simulations have provided insights into the spontaneous polarization caused by band bending and the related carrier distribution, with an optimized EOT 0.312 nm giving a favorable threshold between the depletion and enhancement modes of the MOSFET device.
-Majority carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits.
+The discussion also delves into the influence of spontaneous polarization magnitude, a crucial factor determining the conditions for achieving Ohmic contact and the formation of the channel 2DEG/2DHG.
+The majority of carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits.
 On the basis of the surface polarization strategy adopted here, it might be possible to form a MOSFET from any JTMD, provided that the 2D material possesses a strong intrinsic out-of-plane dipole.

### v3

@@
-Utilizing the Janus MoSSe surface polarization property in complementary metal–oxide–semiconductor field effect transistor design
+Utilizing the Janus MoSSe surface polarization in designing complementary metal-oxide-semiconductor field-effect transistors
 Abstract
 Janus transition metal dichalcogenides (JTMDs) have attracted much attention because of their outstanding electronic and optical properties.
-The additional out-of-plane dipole in JTMDs can form n- and p-like contacts, and this may be used in device applications such as pin diodes and photovoltaic cells.
-In this study, we exploit this property to design n- and p-type metal–oxide–semiconductor field effect transistors (MOSFETs).
-First, we use density-functional theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe.
+The additional out-of-plane dipole in JTMDs can form n- and p-like contacts, and this may be used in device applications, such as pin diodes and photovoltaic cells.
+In this study, we exploit this property to design n- and p-type metal-oxide-semiconductor field-effect transistors (MOSFETs).
+First, we use density-functional-theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe.
 The intrinsic dipole of MoSSe causes band bending at both the metal/SMoSe and SMoSe/metal interfaces, resulting in electron and hole accumulation to form p-type, n-type, and even Ohmic contact regions.
-We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, based on which we design complementary MOSFETs.
+We incorporate this property into a two-dimensional finite-element-based Poisson-drift-diffusion solver to perform simulations, based on which we design complementary MOSFETs.
 Our results demonstrate that JTMDs can be used to make n- and p-MOSFETs in the same layer without the need for any extra doping.
 I Introduction
-In the field of two-dimensional (2D) semiconducting materials, research on transition metal dichalcogenides (TMDs) has been flourishing in recent years owing to their remarkable properties, such as relatively high carrier mobilities,Chhowalla, Jena, and Zhang (2016a); Chen et al.
+In the field of two-dimensional (2D) semiconducting materials, research on transition metal dichalcogenides (TMDs) has been flourishing in recent years owing to their remarkable properties, such as relatively high carrier mobilities,Chhowalla, Jena, and Zhang (2016); Chen et al.
 (2022a); Radisavljevic et al.
-(2011) suitable bandgap energies,Ye, Zhang, and Yap (2017); Chaves et al.
+(2011) suitable band-gap energies,Ye, Zhang, and Yap (2017); Chaves et al.
 (2020) and an atomic thickness that is modifiable.Serna et al.
@@
 (2021); Zhao et al.
-(2021), especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs).
+(2021), especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field-effect transistors (FETs).
 Liu et al.
-(2021); Chhowalla, Jena, and Zhang (2016b); Chen, Chen, and Wu (2022) These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials.
+(2021); Chhowalla, Jena, and Zhang (2016); Chen, Chen, and Wu (2022) These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials.
 For example, MoS2 and WS2 are two commonly selected materials for n-type channels, and -MoTe2 and WSe2 are both good candidates for p-type channels.Radisavljevic et al.
@@
 (2021)
-It is well known that the contact characteristics of a FET device will strongly affect its – characteristics and performance.
-For low contact resistance, Ohmic contacts are usually required at the semiconductor–metal interface to enhance the electrical performance of FETs while maintaining a linear – curve.Sze, Li, and Ng (2021); Allain et al.
+It is well known that the contact characteristics of a FET device will strongly affect its - characteristics and performance.
+For low contact resistance, Ohmic contacts are usually required at the semiconductor-metal interface to enhance the electrical performance of FETs while maintaining a linear - curve.Sze, Li, and Ng (2021); Allain et al.
 (2015); Zheng et al.
@@
 For n-FETs, a material with a smaller work function is needed, whereas a metal with a high work function is often selected for p-FETs.
-Several studies have shown that TMD/metal heterostructures such as MoS2/Cu, WS2/Cu, and WSe2/Ti could provide n-type SB contacts,Kang et al.
-(2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe2/Pd and WSe2/Au.Kang et al.
+Several studies have shown that TMD/metal heterostructures, such as MoS2/Cu, WS2/Cu, and WSe2/Ti could provide n-type SB contacts,Kang et al.
+(2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations, such as WSe2/Pd and WSe2/Au.Kang et al.
 (2014); Kong et al.
 (2020) However, only a specific TMD/metal combination is acceptable for Ohmic contact formation, making TMD-based FET contact design more challenging.Zhao and Schwingenschlögl (2022)
-To combine n- and p-type FETs into a complementary metal–oxide–semiconductor FET (CMOSFET) device, horizontally combing two different TMD channels to form a CMOSFET is a common approach.Liu et al.
+To combine n- and p-type FETs into a complementary metal-oxide-semiconductor FET (CMOSFET) device, horizontally combing two different TMD channels to form a CMOSFET is a common approach.Liu et al.
 (2014); Das, Dubey, and Roelofs (2014); Cho, Park, and Kwon (2015); Xiong et al.
@@
 (2021) Furthermore, two contact metals need to be considered separately, resulting in higher processing complexity and a more time-consuming fabrication procedure.
-Recently, ambipolar TMDs such as MoTe2Lin et al.
+Recently, ambipolar TMDs, such as MoTe2Lin et al.
 (2014) and ReSe2Lee et al.
-(2019) have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
-The use of a novel type of materials, namely, Janus transition metal dichalcogenides (JTMDs), MXY, where M transition metal and X, Y S, Se, or Te, with X Y, is another possible way to fabricate a CMOSFET using a single material.
+(2019), have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
+The use of an alternative type of materials, namely, Janus transition metal dichalcogenides (JTMDs), MXY, where M transition metal and X, Y S, Se, or Te, with X Y, is another possible way to fabricate a CMOSFET using a single material.
 In recent years, several JTMDs have been successfully synthesized via different techniques.
@@
 Lin et al.
-(2020) prepared Janus WSSe by pulsed laser deposition.
+(2020) prepared Janus WSSe by pulsed-laser deposition.
 Previous theoretical studies of these JTMDs have provided unmistakable proof that their phonon spectra are stable and devoid of imaginary frequencies, verifying their fabrication stability.Petrić et al.
 (2021); Chaurasiya, Gupta, and Dixit (2019) Compared with TMDs, JTMDs have two different chalcogen-atom layers, which create an asymmetric sandwich configuration that provides an out-of-plane electric field.
-This out-of-plane field has the characteristic of a dipole, which enables potential applications in piezoelectric, spintronics, catalysts, and gas sensing devices.Zhang et al.
+This out-of-plane field has the characteristic of a dipole, which enables potential applications in piezoelectric, spintronics, catalysts, and gas-sensing devices.Zhang et al.
 (2022, 2020); Kong et al.
@@
 (2020) The remainder of this paper is organized as follows.
-First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, Gr/trilayer AA′A stacking MoSSe/Gr and A′A stacking bilayer MoSSe(2L-MoSSe)/Al (111) face and 2L-MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT).
-Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson–drift–diffusion solver (2D-DDCC) developed in our laboratory.Wu
+First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, Gr/trilayer AA′A stacking MoSSe/Gr and A′A stacking bilayer MoSSe(2L-MoSSe)/Al (111) face and 2L-MoSSe/Cu (111) face structures, via first-principles density-functional theory (DFT).
+Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson-drift-diffusion solver (2D DDCC) developed in our laboratory.Wu
 II Computational details
@@
 (2009) an open-source DFT calculation software package employing the generalized gradient approximation (GGA) of the exchange-correlation functional and norm-conserving VanderbiltHamann, Schlüter, and Chiang (1979) pseudopotentials combined with DFT-D3 van der Waals correction.Grimme et al.
-(2010) An Monkhorst–Pack sampling is used.
+(2010) An Monkhorst-Pack sampling is used.
 The geometry is refined until the maximal residual forces are less than 0.015 eV/Å.
@@
 For layered MoSSe, we construct an AA′A stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here.
-The only difference is that the mesh is set to in the MoSSe band structure calculation.
+The only difference is that the mesh is set to in the MoSSe band-structure calculation.
 II.2 Device simulation
@@
 This solver was developed in our laboratory and has been used on many different kinds of devices.
-More detailed information can be found on our website.Wu Both the Poisson and drift–diffusion equations are solved self-consistently by the solver.
+More detailed information can be found on our website.Wu Both the Poisson and drift-diffusion equations are solved self-consistently by the solver.
 The equations are as follows:
 where is the dielectric constant, is the electric potential, is the acceptor’s doping concentration, is the donor’s doping concentration, is the electron concentration, is the hole concentration, and are the current densities of electrons and holes, respectively, and and are the electron and hole diffusion coefficients, respectively.
-This solver serves as the basis for all of the device simulations discussed in this article, except the MoSSe channel region calculations.
-| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO2 and Al2O3 parameters are given inherently in 2D-DDCC.Wu |
+This solver serves as the basis for all of the device simulations discussed in this paper, except the MoSSe channel region calculations.
+| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO2 and Al2O3 parameters are given inherently in 2D DDCC.Wu |
 | 1.212 eV | Average band gap | |
@@
 Notably, when Gr contacts a three-layer MoSSe material on the Se side (depicted in red), the Fermi level of the entire system is observed to be 0.38 eV lower than the conduction band of the MoSSe material, indicating an n-type Schottky barrier.
-Conversely, when Gr contacts a three-layer MoSSe material on the S side (depicted in green), a p-type Ohmic contact is observed at the Gamma point.
+Conversely, when Gr contacts a three-layer MoSSe material on the S side (depicted in green), a p-type Ohmic contact is observed at the point.
 These findings align with previous studies on JTMD contacts.Palsgaard et al.
@@
 Consequently, the Fermi level becomes higher than the Dirac cone, resulting in intrinsically n-doped Gr.
-Due to the tilting band edge caused by , the energy difference between the conduction band edge and the Fermi level is smaller than the valence band edge offset at the Se/Gr side, suggesting the formation of an n-region.
-On the other hand, the S-side attachment elevates the band energy of the Gr, with the valence band edge intersecting the Fermi level, indicating an intrinsic p-doped Gr and a p-contact region.
+Due to the tilting band edge caused by , the energy difference between the conduction band edge and the Fermi level is smaller than the valence-band edge offset at the Se/Gr side, suggesting the formation of an n region.
+On the other hand, the S-side attachment elevates the band energy of the Gr, with the valence-band edge intersecting the Fermi level, indicating an intrinsic p-doped Gr and a p-contact region.
 Nevertheless, the Dirac cone of both Gr in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al.
@@
 Figures 3(d)–3(g) show the projected band structures, and Figs.
-3(d)–3(f) representing the Se-face contact of MoSSe with the electrode material, and Figs.
-3(e)–3(g) representing the S-face attachment.
+3(d)–3(f) represent the Se-face contact of MoSSe with the electrode material, and Figs.
+3(e)–3(g) represent the S-face attachment.
 For Al (111) face attachment to the Se and S faces of bilayer MoSSe (2L-MoSSe) [Figs.
@@
 Different from typical TMDs, the internal electric field in JTMD plays a crucial role in augmenting the density of interface states.
-Specifically, in MoSSe, this electric field fosters the generation of electronic energy levels at the metal-semiconductor interface via chemisorption-like bonding, leading to an increased density of interface states between the two materials.
-Zhao and Schwingenschlögl (2022) These additional bonding states, resulting from in MoSSe, are delineated in density of states (DOS) plots for all systems, indicated by a dotted frame.Gong et al.
+Specifically, in MoSSe, this electric field fosters the generation of electronic energy levels at the metal-semiconductor interface via chemisorptionlike bonding, leading to an increased density of interface states between the two materialsZhao and Schwingenschlögl (2022).
+These additional bonding states, resulting from in MoSSe, are delineated in density of states (DOS) plots for all systems, indicated by a dotted frame.Gong et al.
 (2010) Due to the additional bonding states caused by chemical adsorption, it becomes challenging to correlate the band edge of MoSSe closest to the metal layer.
-We correlate the band edge distribution of the closest-to-metal layer of MoSSe (1L MoSSe) based on the direction of previously mentioned and the peak present at the Ev edge of MoSSe layers, and depict the corresponding band edge with DOS on the right side of Figs.
+We correlate the band-edge distribution of the closest-to-metal layer of MoSSe (1L MoSSe) based on the direction of previously mentioned and the peak present at the Ev edge of MoSSe layers, and depict the corresponding band edge with DOS on the right side of Figs.
 3(d) and (e).
 When the Se face contacts with Al [Fig.
-3(d)], it can be observed that the Fermi level coincides with the outer MoSSe (2L MoSSe) at the valence band edge.
+3(d)], it can be observed that the Fermi level coincides with the outer MoSSe (2L MoSSe) at the valence-band edge.
 Due to the upward direction of , the energy band of the inner layer of MoSSe (1L MoSSe) shifts towards lower energy.
@@
 (2010); Zhao and Schwingenschlögl (2022) Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111) [Fig.
-3(f) and (g)], indicating Cu can also bring Ohmic contact properties as 2L-MoSSe attches.
-The chemisorption-like property imparts the advantage concerning tunneling barrier reduction at the contact region, thus improving device performance.
+3(f) and (g)], indicating Cu can also bring Ohmic contact properties as 2L-MoSSe attaches.
+The chemisorptionlike property imparts the advantage concerning tunneling barrier reduction at the contact region, thus improving device performance.
 Zhao and Schwingenschlögl (2022); Gong et al.
@@
 An n-MOSFET can be created similarly, as illustrated to the right of the dashed line, with the only difference being that in this case, it is the Se face, rather than the S face, to which metals are attached from n-type regions at the source and drain.
-This horizontal CMOSFET design has two advantages: (1) the Janus MoSSe reduces the complexity of material selection since only one channel material is used to form n- and p-MOSFETs; (2) this structure enables layer-by-layer build-up during the fabrication process, which makes the process more reliable.
-To investigate the electron band characteristics of the designed MoSSe MOSFETs, the 2D-DDCC solver is used to simulate device characteristics.
-We design our device using MoSSe of 5 nm channel thickness to represent MoSSe multilayers since this thickness ensures there is sufficient charge flowing in the channel region.
-All the parameters used in the 2D-DDCC simulation are given in Table 1.
-In our device simulation setting, the metal work function is set at the mid-gap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst-case metal selection scenario.
-Cross-sectional views of the band diagram are shown in Figs.
+This horizontal CMOSFET design has two advantages: (1) the Janus MoSSe reduces the complexity of material selection since only one channel material is used to form n- and p-MOSFETs; (2) this structure enables layer-by-layer buildup during the fabrication process, which makes the process more reliable.
+To investigate the electron-band characteristics of the designed MoSSe MOSFETs, the 2D DDCC solver is used to simulate device characteristics.
+We design our device using MoSSe of 5-nm channel thickness to represent MoSSe multilayers since this thickness ensures there is sufficient charge flowing in the channel region.
+All the parameters used in the 2D DDCC simulation are given in Table 1.
+In our device simulation setting, the metal work function is set at the midgap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst-case metal selection scenario.
+Cross-section views of the band diagram are shown in Figs.
 4(b) and 4(c) for the p- and n-MOSFETs, respectively.
@@
 4(c)] at the interface of the MoSSe region.
-This band edge bending, originating from , plays a pivotal role in facilitating the formation of ohmic contacts within the device, which is in good agreement with the aforementioned DFT calculations.
+This band-edge bending, originating from , plays a pivotal role in facilitating the formation of Ohmic contacts within the device, which is in good agreement with the aforementioned DFT calculations.
 The difference in physical behavior between the different sides of MoSSe attachment corresponds simply to the difference in sign of here.
@@
 In the absence of assistance, the formation of Ohmic contacts falls short, resulting in the Schottky barrier at the interface.
-The – characteristics of several MoSSe MOSFET designs are now investigated.
+The - characteristics of several MoSSe MOSFET designs are now investigated.
 Figure 5(a) shows a single-gate design with gate length m, which is a common dimension adopted in experimental studies of FETs based on 2D materialsXiong et al.
@@
 The simulations for all devices assume an HfO2 gate oxide of thickness nm.
-Figure 5(d) shows the transfer characteristics (–) of the n- and p-MOSFETs from Figs.
+Figure 5(d) shows the transfer characteristics (-) of the n- and p-MOSFETs from Figs.
 5(a)–5(c).
@@
 The key distinction lies in the opposite orientation of in the JTMD layer.
-The typical – curves for the n- and p-MOSFETs indicate that all three devices exhibit satisfactory enhancement-mode MOSFET behavior.
-Among the three devices, the single-gate design [Figure 5(a)] has the most significant sub-threshold swing (SS) and the lowest on-current .
+The typical - curves for the n- and p-MOSFETs indicate that all three devices exhibit satisfactory enhancement-mode MOSFET behavior.
+Among the three devices, the single-gate design [Figure 5(a)] has the most significant subthreshold swing (SS) and the lowest ONcurrent .
 When a bottom gate is added [Figure 5(b)], the SS drops to around 60.5 mV/dec, while increases by an order of magnitude, indicating an improvement in gate control.
@@
 In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the OFF-state current is decreased by about an order, indicating an improvement in .
-All of these results show that the SS and on/off-current characteristics of the designs considered are those of a typical MOSFET device.
+All of these results show that the SS and ON- or OFF-current characteristics of the designs considered are those of a typical MOSFET device.
 The lower SS and higher consistently observed for all the n-MOSFETs compared with the corresponding p-MOSFETs may originate from the slightly higher mobility of electrons compared with holes.
@@
 6(b) and 6(c) for the n- and p-MOSFETs, respectively.
-Here, we define on/off state switching when .
-First, a SiO2 oxide gate of thickness 1 nm is tested, and the – result is shown by the curve with red circles.
-The device is clearly in the on-state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device.
+Here, we define ON or OFFstate switching when .
+First, a SiO2 oxide gate of thickness 1 nm is tested, and the - result is shown by the curve with red circles.
+The device is clearly in the ON state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device.
 To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an Al2O3 gate with nm (EOT 0.42 nm, blue triangles).
@@
 As can be seen from Figs.
-6(b) and 6(c), both of the – curves for this optimized value of intersect the off-state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased.
+6(b) and 6(c), both of the – curves for this optimized value of intersect the OFF state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased.
 These results for HfO2 with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm.
@@
 Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 2.1 V and 1.8 V for n- and p-MOSFETs, respectively.
-All in all, the – plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
+All in all, the - plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
 It is also worth considering how the value of affects this carrier accumulation phenomenon at the channel MoSSe region.
-To reveal the impact of the intrinsic dipole’s magnitude on device performance, we calculate the two-dimensional electron gas (2DEG) in the n-MOSFET device with a 200 nm length channel, shown in [Figure 7(a)], with various assumed values ranging from cm-2 to cm-2.
+To reveal the impact of the intrinsic dipole’s magnitude on device performance, we calculate the two-dimensional electron gas (2DEG) in the n-MOSFET device with a 200-nm length channel, shown in [Figure 7(a)], with various assumed values ranging from to cm-2.
 These values can be regarded as JTMDs with different intrinsic dipole strengths, providing insights into how the magnitude influences the device’s 2DEG.
@@
 To isolate the impact of , the calculations were performed at zero bias, corresponding to the equilibrium state.
-[Figure 7(b)] illustrates the influence of on the 2DEG concentration.
+Figure 7(b) illustrates the influence of on the 2DEG concentration.
 In cases with high EOT (SiO2, nm), 2DEG exhibits a nearly linear increase with increasing .
@@
 This indicates a depletion mode behavior for MoSSe with HfO2 ( nm) due to the higher initial induced 2DEG.
-The influence of extends to the IV characteristics of MOSFETs.
+The influence of extends to the I-V characteristics of MOSFETs.
 To focus on its impact on the linear 2DEG region, we examined devices with values of , , and cm-2, all designed with a 10 nm thick HfO2.
-As shown in Figure 7(c), an increased shifts the transfer IV curve towards more negative voltages.
+As shown in Figure 7(c), an increased shifts the transfer I-V curve towards more negative voltages.
 This shift reflects the need for a stronger gate voltage to deplete the augmented 2DEG in the channel.
-Moreover, the higher carrier concentration resulting from a higher leads to a greater on-current (Ion).
+Moreover, the higher carrier concentration resulting from a higher leads to a greater ON current (ION).
 Additionally, can influence carrier injection at the contacts.
 For strengths of and cm-2 value of , the device also has a larger gm value due to a smaller contact resistance.
-However, if falls below a certain threshold, devices display non-ideal IV curves near saturation (highlighted by the purple dotted box).
+However, if falls below a certain threshold, devices display nonideal I-V curves near saturation (highlighted by the purple dotted box).
 As detailed in Figure 4(c), insufficient compromises Ohmic contacts, resulting in increased resistance from an additional barrier at the contact regions.
@@
 For p-MOSFETs, the same principles apply.
-To understand the distribution of carriers in the device, we export the two-dimensional distribution map of the majority carrier of the device under different operating conditions in Figure 8, which demonstrates the on/off characteristics of the p- and n-MOSFET devices with HfO2 gate with nm.
+To understand the distribution of carriers in the device, we export the two-dimensional distribution map of the majority carrier of the device under different operating conditions in Figure 8, which demonstrates the ONorOFF characteristics of the p- and n-MOSFET devices with HfO2 gate with nm.
 The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm-3 (green color).
 The carrier concentration gradually decays as the position moves away from the contact region.
-Owing to the 10 nm gate metal isolation design, two bump-like regions of high carrier concentration appear next to the source and drain areas.
+Owing to the 10-nm gate metal isolation design, two bumplike regions of high carrier concentration appear next to the source and drain areas.
 The thicker oxide eliminates the dipole effect of MoSSe, and thus carrier gas remains.
@@
 IV Conclusions
-We have introduced a novel CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe.
+We have introduced an alternative CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe.
 Our investigation has encompassed diverse heterostructures, including Gr/AA′A-MoSSe/Gr and 2L-MoSSe/Al (111) face and 2L-MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT.
-We have also examined the interactions between MoSSe, Gr, and Al and Cu electrodes, identifying characteristics such as Schottky barriers and Ohmic contacts.
+We have also examined the interactions between MoSSe, Gr, and Al and Cu electrodes, identifying characteristics, such as Schottky barriers and Ohmic contacts.
 Our design leverages the strong dipole field of MoSSe to create n- and p-type contact regions, thereby forming a CMOSFET with simplified material selection.
@@
 Device simulations have provided insights into the spontaneous polarization caused by band bending and the related carrier distribution, with an optimized EOT 0.312 nm giving a favorable threshold between the depletion and enhancement modes of the MOSFET device.
-The discussion also delves into the influence of spontaneous polarization magnitude, a crucial factor determining the conditions for achieving Ohmic contact and the formation of the channel 2DEG/2DHG.
+The discussion also delves into the influence of spontaneous polarization magnitude, a crucial factor determining the conditions for achieving Ohmic contact and the formation of the channel 2DEG and 2DHG.
 The majority of carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits.
@@
 Acknowledgements.
-This work was supported by the National Science and Technology Council under Grant Nos. NSTC 112-2221-E-002-215-MY3, NSTC 112-2923-E-002-002, and 12-2221-E-002-214-MY3.AUTHOR DECLARATIONS
-Conflict of Interest
-The authors have no conflicts to disclose
-Author Contributions
+This work was supported by the National Science and Technology Council under Grants No.
+NSTC 112-2221-E-002-215-MY3, No.
+NSTC 112-2923-E-002-002, and No.
+112-2221-E-002-214-MY3.
 Yun-Pin Chiu: conceptualization (equal); data curation (lead); formal analysis (equal); methodology (equal); validation (lead); visualization(lead); writing – original draft (lead).
 Hsin-Wen Huang: conceptualization (equal); data curation (support); formal analysis (equal); methodology (equal); validation (support).
-Yuh-Renn Wu: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Software(lead); Supervision (lead); Writing – review and editing (lead).
-Data Availability
-The data that support the findings of this study are available from the corresponding author upon reasonable request.
+Yuh-Renn Wu conceptualization (equal); formal analysis (equal); funding acquisition (lead); project administration (lead); software(lead); supervision (lead); writing-–review and editing (lead).
+The authors have no conflicts to discloseReferences